# Semidiscrete optimal transport with unknown costs

### v3

@@ 1 Introduction
The optimal transport literature universally assumes that the cost function $c$ is known. In fact, most papers only work with some particular form for $c$, most commonly a Euclidean distance or $p$-norm (Santambrogio, 2015). This paper considers a situation where $c$ is unknown, but can be estimated in a sequential manner from data. Our motivation for this setting is a class of applications where $k\in\left\{1,...,K\right\}$ represents a “choice” or “alternative” available to a decision-maker. One observes a sample of $X$ from the density $f$, then selects an alternative $k$ and incurs a cost $c\left(X,k\right)$. The goal is to assign $X$ values to alternatives in a way that minimizes expected cost while obeying the constraints (2) on how often each alternative can be chosen.
@@
In this paper, we bridge optimal transport theory and optimal learning to formalize a notion of correct selection, and we create a sampling algorithm that makes correct selections at an optimal rate. In the process, we also obtain new results for the more standard setting of semidiscrete optimal transport with known costs. Although the known-cost setting is not the main focus of this research, the results are of some interest, and they lay the groundwork for the analytical technique that we eventually use to handle unknown costs. It is therefore necessary to discuss both settings for a clear understanding of how our analysis proceeds.
1.1 Results for known costs
@@
1.2 Results for unknown costs
@@
2 Optimal policies in semidiscrete optimal transport
We begin by discussing the optimal solution to (1)-(4) when $c$ is known. This solution will provide us with a performance metric for the problem that we will eventually tackle. In contextual ranking and selection (Hong et al., 2021), the main metric used to evaluate and compare algorithms is “correct selection,” which occurs when we select the alternative with the lowest cost given $X$. This is what would happen if we removed the probabilistic target constraints (2), because then (1) would be minimized simply by setting $Y=\arg\min_{k}c\left(X,k\right)$. We require an analogous notion of correct selection that would be meaningful in the presence of (2).
It turns out that (1)-(4) is solved by a policy, or decision rule for assigning any observed $X$ to a $k$ value. In other words, $Y$ should be set equal to a certain deterministic function of $X$. The optimal transport literature has characterized this function near-explicitly using Kantorovich duality (Theorem 1.3 in Villani, 2021). To describe and explain its structure, we sketch the key steps in the derivation. These results are well-known and can be found in, e.g., Carlsson et al. (2016) or Genevay et al. (2016), but we summarize them here to provide background.
For notational convenience, we write $c_{k}\left(x\right)=c\left(x,k\right)$, to visually distinguish between costs $c_{k}$ corresponding to different “alternatives,” from which we will eventually make a selection. We do not impose any parametric structure on $c_{k}$. By Kantorovich duality, the problem (1)-(4) has the same optimal value as the functional optimization problem
@@ 2 Optimal policies in semidiscrete optimal transport
[…]
@@
3 Algorithms and main results
Under both known and unknown costs, our goal is to learn $\pi^{*}$ iteratively based on a sequence $\left\{X^{n}\right\}^{\infty}_{n=1}$ of i.i.d. samples from the data-generating density $f$. Although our main focus is on unknown costs, the analysis builds on new results that hold under known costs. For this reason, our presentation separates the two cases. Section 3.1 states our assumptions and main results for the known-cost setting, and Section 3.2 does the same for unknown costs.
3.1 Algorithm and main results: known costs
@@
where $\left\{\alpha_{n}\right\}^{\infty}_{n=0}$ is a stepsize sequence satisfying the usual conditions $\sum_{n}\alpha_{n}=\infty$, $\sum_{n}\alpha^{2}_{n}<\infty$. The gradient is bounded, so the convergence $g^{n}\rightarrow g^{*}$ easily follows from classical SA theory (Kushner & Yin, 2003). Note that (12) can be applied without knowing the density $f$ from which each $X^{n+1}$ is sampled. This is an attractive feature of SA, as data-generating distributions are extremely difficult to estimate. If the goal is to maximize expected costs, as in the motivating example from Section 1, we can use (12) with a trivial sign change.
@@
From (11), we see that adding a constant to every component of $g^{*}$ does not change the optimal decision. Therefore, we only need to learn the differences $g^{*}_{k}-g^{*}_{1}$ for $k>1$, so we may assume without loss of generality that $g^{*}_{1}=0$. We impose a simple regularity condition ensuring that $F\left(g,X\right)$ has a unique argmin for any choice of $g$.
@@
For any $1\leq j,k\leq K$ with $j\neq k$, the random variable $c_{j}\left(X\right)-c_{k}\left(X\right)$ is absolutely continuous with respect to Lebesgue measure.
The next set of assumptions concerns the “richness” of the data-generating distribution. It is based on the recent optimal learning literature, which has found, in various settings, that desirable convergence rates require the observed values of $X$ to exhibit a sufficient degree of random variation. This is made more precise as follows.
Assumption 3.2.
Suppose that there exist constants $\kappa_{1},...,\kappa_{4}>0$ (which may depend on $g^{*}$) such that:
@@
    has a density that is bounded within $\left[\kappa_{2},\kappa_{3}\right]$ on the interval $\left[-\kappa_{1},\kappa_{1}\right]$.
@@
    The support of $X$ is bounded, i.e., $\|X\|_{2}\leq\kappa_{4}$.
These assumptions are fairly typical for recent work on optimal learning. Assumption 3.2(i) is very similar to the margin conditions in Assumption A2 of Goldenshluger & Zeevi (2013), Assumption 2 of Bastani & Bayati (2020), and Assumption 2 of Bastani et al. (2021). Essentially, we require sufficient random variation near the decision boundary $\left\{Z_{j,k}=0\right\}$, as this is where we have the greatest difficulty identifying whether $j$ or $k$ is better. Assumption 3.2(ii) is also made by the above-cited papers and many others.
The algorithm is given by (12) using a simple stepsize $\alpha_{n}=\frac{\alpha}{n+1}$ with $\alpha>0$ being a pre-specified constant. Our main result is a rate bound on the convergence in $L^{2}$ of the sequence $\left\{\delta^{n}\right\}^{\infty}_{n=0}$, with $\delta^{n}=g^{n}-g^{*}$.
@@
Let Assumptions 3.1-3.2 hold. There exists some $\bar{\alpha}>0$, which depends only on $\kappa_{1}$ and $\kappa_{2}$, such that $\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)\leq\frac{D}{n}$ if $\alpha>\bar{\alpha}$.
Thus, under Assumptions 3.1-3.2, classical SA achieves the canonical $\mathcal{O}\left(\frac{1}{n}\right)$ rate. It is interesting to note that this does not happen in general: one usually resorts to Polyak averaging (Polyak & Juditsky, 1992) to recover the optimal convergence rate, which is indeed done in Bach & Moulines (2011). However, in the specific setting of semidiscrete optimal transport with known costs, neither averaging nor strong concavity are necessary as long as the data-generating process is rich in the sense of Assumption 3.2. This result is completely new to our paper, and is valuable for preserving one of the most attractive characteristics of the SA approach to semidiscrete optimal transport, namely its ability to learn (in the limit) the exact optimal policy without resorting to approximation or discretization of the problem.
3.2 Algorithm and main results: unknown costs
@@
Initialize $n=0$, $g^{0}$ and $\hat{\beta}^{0}_{k}$ for $k=1,...,K$. Choose a deterministic stepsize sequence $\left\{\alpha_{n}\right\}^{\infty}_{n=0}$.
@@
Formally, we first define $\hat{\pi}^{n+1}=\arg\min_{j}\hat{c}^{n}_{j}\left(X^{n+1}\right)-g^{n}_{j}$ to be the plug-in estimate of the optimal decision, based on estimates $\hat{\beta}^{n}$ and $g^{n}$. We also define, for each $k=1,...,K$, the set
@@
The estimate $g^{n}_{j}$ is updated using (13), and the regression coefficients are computed using a suitable generalized linear model.
@@
As in Section 3.1, we let $\alpha_{n}=\frac{\alpha}{n+1}$ and $\delta^{n}=g^{n}-g^{*}$. We also let $\Delta^{n}_{k}=\hat{\beta}^{n}_{k}-\beta_{k}$. Recall the distinction between $\hat{\pi}^{n}$, which is the plug-in estimate of the optimal decision, and $\pi^{n}$, which is the semi-myopic policy combining that estimate with forced exploration as defined in (14).
@@
Let Assumptions 3.1-3.2 and 3.3 hold, and suppose that the semi-myopic policy $\pi^{n}$ is used to sample cost functions. There exist $\bar{\alpha},C,C^{\prime}>0$, which depend only on $\kappa_{1},...,\kappa_{6}$, such that
$\mathbb{E}\left(\max_{j}\|\Delta^{n}_{j}\|^{2}_{2}\right)\leq\frac{C}{n},\qquad\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)\leq\frac{C^{\prime}}{n},$
if $\alpha>\bar{\alpha}$.
@@
4 Analysis for known costs
In this section, we return to the setting of known costs and prove Theorem 3.1. As was discussed previously, the main difficulty in the proof is the lack of strong concavity in the objective function. Thus, a crucial first step is the derivation (in Section 4.1) of a kind of “local strong concavity” property that is weaker than strong concavity, but still sufficient to prove (in Section 4.2) the main rate result. We will later use this property in our analysis of unknown costs, but it holds more generally and does not require $c_{k}$ to have any parametric structure. Thus, we are the first to show that semidiscrete optimal transport can be solved at an optimal rate without the need for discretization or smoothing, as long as the data-generating distribution is sufficiently rich.
@@ 4.1 Properties of the objective
We establish two technical results concerning the smoothness of the objective $\mathbb{E}\left(F\left(g,X\right)\right)$. The proof techniques will be used again later on, in Section 5, where these properties will be generalized to the setting of unknown cost functions.
@@
Let Assumptions 3.1-3.2 hold, and take any $\delta\in\mathds{R}_{K}$ satisfying $\delta_{1}=0$. Then,
$\mathbb{E}\left(F\left(g^{*},X\right)\right)-\mathbb{E}\left(F\left(g^{*}+\delta,X\right)\right)-\delta^{\top}p\geq\frac{\kappa_{2}}{2K}\left(\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\}\right).$
@@
Clearly, $\mathbb{E}\left(T_{1}\right)=\delta^{\top}p$. We will work toward a lower bound on $\mathbb{E}\left(T_{2}\right)$, which will require some algebra on (17). Because the indices maximizing $F\left(g+\delta,X\right)$ and $F\left(g,X\right)$ are unique by assumption, we can write

### v4

@@ 1 Introduction
The optimal transport literature universally assumes that the cost function $c$ is known: in fact, most papers only work with some particular form for $c$, most commonly a Euclidean distance or $p$-norm (Santambrogio, 2015). Our paper is the first to focus on a situation where $c$ is unknown, but can be estimated in a sequential manner from data. Our motivation for this setting is a class of applications where $k\in\left\{1,...,K\right\}$ represents a “choice” or “alternative” available to a decision-maker. One observes a sample of $X$ from the density $f$, then selects an alternative $k$ and incurs a cost $c\left(X,k\right)$. The goal is to assign $X$ values to alternatives in a way that minimizes expected cost while obeying the constraints (2) on how often each alternative can be chosen.
@@
In this paper, we bridge optimal transport theory and optimal learning to formalize a notion of correct selection, and we create a sampling algorithm that makes correct selections at an optimal rate. As a byproduct, we also obtain new results for the more standard setting of semidiscrete optimal transport with known costs. Although the known-cost setting is not the main focus of this research, we use it as a starting point to develop the analytical technique that we eventually use to handle unknown costs. We first explain the essential elements of this technique in Section 1.1 below, and then discuss the main results in Section 1.2.
1.1 Stochastic approximation and local strong concavity
@@
1.2 Main contribution: semi-myopic learning of unknown costs
@@
1.3 Organization of the paper
Section 2 presents preliminaries on the semidiscrete optimal transport problem, most importantly the structure of the optimal policy. We then state the algorithm and key theoretical results: Section 3 presents stochastic approximation (Section 3.1) and semi-myopic learning (Section 3.2) together with our main assumptions and rate results. Then, Section 4 presents the main steps of our technical analysis for SA (local strong concavity under known costs), while Section 5 does the same for semi-myopic learning of unknown costs. Section 6 presents numerical examples and insights, and Section 7 concludes.
2 Semidiscrete optimal transport: preliminaries
We begin by reviewing the structure of the semidiscrete optimal transport problem. The literature has characterized the structure of the optimal solution to (1)-(4) when $c$ is known; these results can be found in, e.g., Carlsson et al. (2016) or Genevay et al. (2016), but we summarize them here to provide background, sketching the key steps in the derivation. In brief, the problem (1)-(4) is solved by a policy, or decision rule for assigning any observed $X$ to a $k$ value. In other words, $Y$ should be set equal to a certain deterministic function of $X$. It is necessary to understand this function before we can formally define a performance metric for the problem that we will eventually tackle.
For notational convenience, we write $c_{k}\left(x\right)=c\left(x,k\right)$, to visually distinguish between costs $c_{k}$ corresponding to different “alternatives,” from which we will eventually make a selection. We do not impose any parametric structure on $c_{k}$. By Kantorovich duality (Theorem 1.3 in Villani, 2021), the problem (1)-(4) has the same optimal value as the functional optimization problem
@@ 2 Semidiscrete optimal transport: preliminaries
$\phi\left(X\right)+\psi\left(Y\right)\leq\sum_{k}c_{k}\left(X\right)1_{\left\{Y=k\right\}}\quad\text{a.s.}$ (6)
@@
3 Algorithm and main results
Our overarching goal is to learn $\pi^{*}$ iteratively based on a sequence $\left\{X^{n}\right\}^{\infty}_{n=1}$ of i.i.d. samples from the data-generating density $f$. Ultimately, we wish to do this in a setting where the cost functions $c_{k}$ are unknown and have to be estimated. Our algorithmic approach combines two elements: stochastic approximation (for learning the optimal dual variables $g^{*}$) and semi-myopic exploration (for learning the costs).
Integrating these elements is quite challenging. Therefore, we begin by examining stochastic approximation separately, in a setting where the costs are known. Section 3.1 states the assumptions and main results needed to establish convergence rates for SA in semidiscrete optimal transport. These results, particularly the local strong concavity property, will continue to play a vital role when we switch to the unknown-cost setting. Section 3.2 presents our algorithm and main results for that case.
3.1 Stochastic approximation: known costs
@@
where $\left\{\alpha_{n}\right\}^{\infty}_{n=0}$ is a stepsize sequence satisfying the usual conditions $\sum_{n}\alpha_{n}=\infty$, $\sum_{n}\alpha^{2}_{n}<\infty$. The gradient is bounded, so the convergence $g^{n}\rightarrow g^{*}$ easily follows from classical SA theory (Kushner & Yin, 2003). The only input to (12) is the stepsize $\alpha_{n}$. The algorithm requires knowledge of the targets $p_{k}$ and the cost functions $c_{j}$, but does not require us to know the density $f$ from which each $X^{n+1}$ is sampled. This is an attractive feature of SA, as data-generating distributions are extremely difficult to estimate. If the goal is to maximize expected costs, as in the motivating example from Section 1, we can use (12) with a trivial sign change.
@@
From (11), we see that adding a constant to every component of $g^{*}$ does not change the optimal decision. Therefore, we only need to learn the differences $g^{*}_{k}-g^{*}_{1}$ for $k>1$, so we may assume without loss of generality that $g^{*}_{1}=0$. Our main assumption concerns the “richness” of the data-generating distribution. It is based on the recent optimal learning literature, which has found, in various settings, that desirable convergence rates require the observed values of $X$ to exhibit a sufficient degree of random variation. This is made more precise as follows.
@@
Let $Z$ be the random vector whose $j$th component is equal to $c_{j}\left(X\right)-g^{*}_{j}$. Suppose that there exist constants $\kappa_{1},...,\kappa_{4}>0$ (which may depend on $g^{*}$) such that:
  1. i)
    The density of $Z$ exists and is bounded below by $\kappa_{2}$ on $\left[-\kappa_{1},\kappa_{1}\right]^{K}$, and bounded above by $\kappa_{3}$ everywhere on its support.
  2. ii)
    Both $X$ and $Z$ have bounded support, i.e., $\|X\|_{2}\leq\kappa_{4}$ and $\|Z\|_{2}\leq\kappa_{4}$.
Assumption 3.1(i) is a so-called “margin condition,” requiring sufficient random variation near the decision boundary (where multiple components of $Z$ are close to zero). Conditions of this type in the existing literature include Assumption A2 of Goldenshluger & Zeevi (2013), Assumption 2 of Bastani & Bayati (2020), and Assumption 2 of Bastani et al. (2021), all of which also assume boundedness as in Assumption 3.1(ii). The main difference between these papers and our work is that they impose margin conditions for pairs of choices (in our setting, these would be the pairwise differences $Z_{j}-Z_{k}$), while we make the slightly stronger assumption that the margin condition holds on the entire multivariate density of $Z$. The distinctions between assumptions are due to the fact that, in the above-cited papers, the penalties $g^{*}_{j}$ are not present at all, and the ultimate goal is only to learn the cost functions, which is done in a “marginal” way, by observing one $k$ at a time. In our setting, even when the costs are known, we still have the problem of learning $g^{*}_{j}$, which is done “jointly” by updating our estimates of all the penalties in each iteration of (12). It therefore becomes necessary to consider the decision boundary in $K$ dimensions.
When Assumption 3.1 is imposed, one then obtains certain useful technical properties for the pairwise differences. These properties will be useful for our analysis later. We state them below; the proof is deferred to the Appendix. Note that the second property considers the pairwise differences jointly by minimizing over them, which is one reason why Assumption 3.1 holds on the multivariate density of $Z$.
Lemma 3.1.
Let Assumption 3.1 hold. There exist constants $\kappa^{\prime}_{1},...,\kappa^{\prime}_{4}>0$ (which depend on $\kappa_{1},...,\kappa_{4}$) such that:
@@
    has a density that is bounded below by $\kappa^{\prime}_{2}$ on the interval $\left[-\kappa^{\prime}_{1},\kappa^{\prime}_{1}\right]$, and bounded above by $\kappa^{\prime}_{3}$ everywhere.
@@
    For any $j\neq k$ and any $0<\kappa<\kappa^{\prime}_{1}$, we have
    $P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid 0<Z_{j,k}\leq\kappa\right)\geq\kappa^{\prime}_{4}.$
Under known costs, our algorithm for learning $g^{*}$ is given by (12) using a simple stepsize $\alpha_{n}=\frac{\alpha}{n+1}$ with $\alpha>0$ being a pre-specified constant. Our main result is a rate bound on the convergence in $L^{2}$ of the sequence $\left\{\delta^{n}\right\}^{\infty}_{n=0}$, with $\delta^{n}=g^{n}-g^{*}$.
@@
Let Assumption 3.1 hold. There exists some $\bar{\alpha}>0$, which depends only on $\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$, such that $\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)\leq\frac{D}{n}$ if $\alpha>\bar{\alpha}$.
Thus, under Assumption 3.1, classical SA achieves the canonical $\mathcal{O}\left(\frac{1}{n}\right)$ rate. It is interesting to note that this does not happen in general: one usually resorts to Polyak averaging (Polyak & Juditsky, 1992) to recover the optimal convergence rate, which is indeed done in Bach & Moulines (2011). However, in the specific setting of semidiscrete optimal transport with known costs, neither averaging nor strong concavity are necessary as long as the data-generating process is rich in the sense of Assumption 3.1. This result is completely new to our paper, and is valuable for preserving one of the most attractive characteristics of the SA approach to semidiscrete optimal transport, namely its ability to learn (in the limit) the exact optimal policy without resorting to approximation or discretization of the problem.
The result of Theorem 3.1 depends on two parameters, the threshold $\bar{\alpha}$ used by the stepsize, and the constant $D$ in the bound. We briefly clarify their dependence on the constants $\kappa_{1},...,\kappa_{4}$ obtained from Assumption 3.1. The threshold $\bar{\alpha}$ is the main source of this dependence, because $D$ is essentially an increasing function of $\alpha>\bar{\alpha}$. Our proofs indicate that $\bar{\alpha}$ is decreasing in $\kappa_{1}$ and $\kappa_{2}$ (that is, $\bar{\alpha}$ is larger when these constants are smaller) and increasing in $\kappa_{3}$ and $\kappa_{4}$. It is not possible to precisely quantify each relationship because, for example, $\bar{\alpha}$ depends on the support of $Z$ (which is controlled by $\kappa_{4}$) indirectly, through the support of the pairwise differences $Z_{j,k}$. However, roughly speaking, expanding the support of $X$ generally increases the support of $Z_{j,k}$ and thus the threshold.
3.2 Semi-myopic learning: unknown costs
@@
Initialize $n=0$, $g^{0}$ and $\hat{\beta}^{0}_{k}$ for k=1,,Kk=1,...,K. Choose a deterministic stepsize sequence $\left\{\alpha_{n}\right\}^{\infty}_{n=0}$.
@@
Formally, we first define $\hat{\pi}^{n+1}=\arg\min_{j}\hat{c}^{n}_{j}\left(X^{n+1}\right)-g^{n}_{j}$ to be the plug-in estimate of the optimal decision, based on estimates $\hat{\beta}^{n}$ and $g^{n}$. We also define, for each k=1,,Kk=1,...,K, the set
@@
The estimate $g^{n}_{j}$ is updated using (13), and the regression coefficients are computed using a suitable generalized linear model. Thus, in order to implement the algorithm, we require only an exploration parameter $a>0$ and a stepsize sequence; as in Section 3.1, we will use $\alpha_{n}=\frac{\alpha}{n+1}$, so only a single parameter $\alpha$ needs to be specified.
@@
As before, we let $\delta^{n}=g^{n}-g^{*}$. We also define $\Delta^{n}_{k}=\hat{\beta}^{n}_{k}-\beta_{k}$. Recall the distinction between $\hat{\pi}^{n}$, which is the plug-in estimate of the optimal decision, and $\pi^{n}$, which is the semi-myopic policy combining that estimate with forced exploration as defined in (14).
@@
Let Assumptions 3.1 and 3.2 hold, and suppose that the semi-myopic policy $\pi^{n}$ is used to sample cost functions. There exists $\bar{\alpha}$ depending only on $\kappa_{1},...,\kappa_{6}$, such that, when $\alpha>\bar{\alpha}$, we have
$\mathbb{E}\left(\max_{j}\|\Delta^{n}_{j}\|^{2}_{2}\right)\leq\frac{C}{n},\qquad\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)\leq\frac{C^{\prime}}{n}$
for some $C,C^{\prime}>0$ depending on $\alpha,\kappa_{1},...,\kappa_{6}$.
As in the case of Theorem 3.1, we may briefly discuss the dependence of the threshold $\bar{\alpha}$ on the constants $\kappa_{1},...,\kappa_{6}$. As before, the threshold is decreasing in $\kappa_{1},\kappa_{2}$ and increasing in $\kappa_{3}$. The dependence on $\kappa_{4}$ is more complex because the support of $X$ indirectly influences various constants in the proof, but it is possible to say that the threshold becomes increasing in $\kappa_{4}$ for large enough values of this constant. The new constants $\kappa_{5},\kappa_{6}$ that are specific to the unknown-cost setting do not influence $\bar{\alpha}$, only the rate constants $C,C^{\prime}$. Unfortunately, the dependence of $C,C^{\prime}$ on $\kappa_{1},...,\kappa_{6}$ is non-monotonic and very difficult to characterize.
@@
4 Stochastic approximation: technical analysis
Earlier, in Section 3, we separated the presentation of the main results according to the core components of our algorithm: first, stochastic approximation under known costs (Theorem 3.1), and then semi-myopic learning under unknown costs (Theorems 3.2-3.3). We make the same distinction when presenting the proofs. In this section, we confine ourselves to the setting of Section 3.2 and prove Theorem 3.1. The unknown-cost case will be treated later in Section 5.
As was discussed previously, the main technical challenge in establishing Theorem 3.1 is the lack of strong concavity in the objective function. A crucial first step is the derivation (in Section 4.1) of a kind of “local strong concavity” property that is weaker than strong concavity, but still sufficient to prove (in Section 4.2) the main rate result. We will later use this property in our analysis of unknown costs, but it holds more generally and does not require $c_{k}$ to have any parametric structure. For that reason, the results in this section are of some stand-alone interest, because we are the first to show that semidiscrete optimal transport can be solved at an optimal rate without the need for discretization or smoothing, as long as the data-generating distribution is sufficiently rich.
@@ 4.1 Properties of the objective
We establish two technical results concerning the smoothness of the objective $\mathbb{E}\left(F\left(g,X\right)\right)$. The proof techniques will be used again later on, in Section 5, where these properties will be generalized to the setting of unknown cost functions. We will assume Assumption 3.1 throughout, denoting by $\kappa^{\prime}_{1},...,\kappa^{\prime}_{4}$ the constants obtained from Lemma 3.1.
@@
Let Assumption 3.1 hold, and take any $\delta\in\mathds{R}^{K}$ satisfying $\delta_{1}=0$. Then,
$\mathbb{E}\left(F\left(g^{*},X\right)\right)-\mathbb{E}\left(F\left(g^{*}+\delta,X\right)\right)-\delta^{\top}p\geq\frac{\kappa^{\prime}_{2}\kappa^{\prime}_{4}}{4K}\left(\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}\right).$
@@
Clearly, $\mathbb{E}\left(T_{1}\right)=\delta^{\top}p$. We will work toward a lower bound on $\mathbb{E}\left(T_{2}\right)$, which will require some algebra on (17). Because the indices minimizing $F\left(g+\delta,X\right)$ and $F\left(g,X\right)$ are unique by Assumption 3.1, we can write